### v1

# Not All Attention is Needed: Parameter and Computation Efficient Transfer Learning for Multi-modal Large Language Models

Abstract
In this paper, we propose a novel parameter and computation efficient tuning method for Multi-modal Large Language Models (MLLMs), termed Efficient Attention Skipping (EAS). Concretely, we first reveal that multi-head attentions (MHAs), the main computational overhead of MLLMs, are often redundant to downstream tasks. Based on this observation, EAS evaluates the attention redundancy and skips the less important MHAs to speed up inference. Besides, we also propose a novel propagation-of-information adapter (PIA) to serve the attention skipping of EAS and keep parameter efficiency, which can be further re-parameterized into feed-forward networks (FFNs) for zero-extra latency. To validate EAS, we apply it to a recently proposed MLLM called LaVIN and a classic VL pre-trained model called METER, and conduct extensive experiments on a set of benchmarks. The experiments show that EAS not only retains high performance and parameter efficiency, but also greatly speeds up inference speed. For instance, LaVIN-EAS can obtain 89.98% accuracy on ScineceQA while speeding up inference by 2.2 times to LaVIN222The source code is anonymously released at: https://anonymous.4open.science/r/EAS-C0E5..
1 Introduction
Recently, the great success of Large Language Models (LLMs) also sparks an influx of interest in extending these giant model to more modalities, i.e., Multi-modal Large Language Models (MLLMs) (Luo et al., 2023b; Liu et al., 2023a; Li et al., 2023). However, compared with the unimodal LLMs, MLLMs inevitably involve much more computation with the introduction of new modalities (Bao et al., 2022; Alayrac et al., 2022; Li et al., 2023; Liu et al., 2023a), and their adaptions to downstream tasks also take more training and parameter expenditure (Kim et al., 2021b; Li et al., 2021, 2022). In this case, numerous efforts have been recently devoted to the efficient tuning of MLLMs (Zhang et al., 2023; Gao et al., 2023b; Wu et al., 2023b).
Among these advancements, Wu et al. (2023b) propose a non-trivial task for large-scaled pre-trained models, termed parameter and computation efficient transfer learning (PCETL). A key intuition behind PCETL is that the sheer size of model parameters and computation is critical for large-scale pre-training but redundant to specific tasks. In this case, for downstream task adaptions, PCETL not only requires to reduce the number of updated parameters, akin to PETL (Houlsby et al., 2019; Hu et al., 2022; Sung et al., 2022b), but also needs to remove the redundant modules for better efficiency. Wu et al. (2023b) also propose a strong baseline called dynamic architecture skipping (DAS) to evaluate and skip the redundant Transformer layers with lightweight adapters, thereby saving both parameter and computation overhead on downstream tasks.
Despite the effectiveness, we think that this strong baseline still leaves ample room to improve for MLLMs. Above all, DAS (Wu et al., 2023b) only considers the entire Transformer layer for redundancy evaluation, and a granular scheme deserves more explorations. As shown in Fig. 1-(a), multi-head attention (MHA) consumes 35% more time than feed-forward network (FFN), while its parameter size is only half. From existing literature (Martins et al., 2020; Tolstikhin et al., 2021), we also notice that MHA is mainly in charge of dependency modeling, and FFN is often used to improve model capacity. As shown in Fig. 1-(b), skipping a certain number of MHAs does not affect performance, while it will results in a drastic drop with FFN. In this case, we assume that FFN stores more pre-training knowledge for downstream task adaption. In contrast, for a simple task, the repeated attention modeling may be of limited significance. Based on these observations, we make an argument in this paper: “Not all attention is needed for the parameter and computation efficient tuning of MLLMs.”
Another shortcoming of DAS is its skip connections with adapters, which still incurs non-negligible latency during inference. As revealed in recent progresses (Luo et al., 2023b; Gao et al., 2023b), the additional branches added by adapters inevitably slow down inference even with very limited FLOPs (Luo et al., 2023b). The same as shown in Fig.1, the latency caused by adapters takes up to 37.9% inference time of LaVIN (Luo et al., 2023b). This drawback somewhat undermines DAS’s efficiency for MLLMs.
To overcome these limitations, we propose a novel approach for PCETL of MLLMs in this paper, termed efficient attention skipping (EAS). The principle of EAS is to only skip the redundant attention modeling for better efficiency and performance. To avoid the latency caused by adapters, we also equip EAS with an innovative parameter efficient module called propagation-of-information adapter (PIA), which can not only replace the skipped MHAs for adaption, but also can be seamlessly re-parameterized into FFNs for zero extra cost during inference. With these innovative designs, EAS can help MLLMs achieve better PCETL on vision-language tasks, i.e., true model acceleration with a small number of trainable parameters.
To validate EAS, we first apply it to a recent MLLM called LaVIN (Luo et al., 2023b), and conduct extensive experiments on the open-set multi-modal question answer benchmark, namely ScienceQA (Lu et al., 2022). To align DAS (Wu et al., 2023b), we also apply EAS to a representative VL pre-trained model called METER (Dou et al., 2022) on three VL benchmarks, i.e., VQA2.0 (Goyal et al., 2017a), NLVR (Suhr et al., 2019a) and Flickr30K (Plummer et al., 2017). The experimental results show that EAS can not only retain high performance and parameter efficiency against existing PETL and PCETL methods (Zhang et al., 2023; Luo et al., 2023b; Wu et al., 2023b), but also boost the inference speed on downstream tasks. For instance, EAS can improve the inference speed by to DAS (Wu et al., 2023b) and to the default LaVIN (Luo et al., 2023b) without performance degradations.
Overall, our contributions can be summarized as follows:
We propose a novel parameter and computation efficient tuning method for MLLMs, termed efficient attention skipping (EAS), which retains the high performance of MLLMs and reduces both parameter and computation expenditures on downstream tasks.
We propose a novel propagation-of-information adapter (PIA) that can be used to serve attention skipping and be fully re-parameterized into MLLMs for true model acceleration.
We apply our EAS to a recent MLLM and a representative VLP model, i.e., LaVIN and METER, on four VL benchmarks. The experiments show the obvious merits of EAS in both parameter and computation efficiencies, e.g., speeding up LaVIN by 2.18 without performance degradations.
2 Related Work
2.1 Vision-language pre-trained models
Recent years have witnessed the rapid development of vision-language pre-trained (VLP) models (Li et al., 2019; Lu et al., 2019; Su et al., 2019; Cho et al., 2021; Wang et al., 2022) on various vision-language (VL) tasks (Goyal et al., 2017b; Suhr et al., 2019b; Lu et al., 2022). These VLP models often adopt pre-trained vision and language models as the encoders, and build a deep fusion branch for cross-modal interactions. The merits of this design are that vision-language alignment can be effectively modeled in the additional fusion branch based on the well learned features of two modalities (Lu et al., 2019; Tan & Bansal, 2019; Li et al., 2021; Dou et al., 2022; Bao et al., 2022).
With the advent of large language models (LLMs) (Zhang et al., 2022; Touvron et al., 2023; OpenAI, 2023), recent researches (Luo et al., 2023b; Liu et al., 2023a; Li et al., 2023) extend these LLMs to vision-language tasks, i.e., multi-modal large language models (MLLMs), where LLMs are directly used for language modeling and multi-modal fusion. More recently, there are attempts of parameter-efficiently transferring LLM to vision-language tasks (Goyal et al., 2017b; Lu et al., 2022). These methods insert lightweight modules like adapter (Zhang et al., 2023) for downstream adaption instead of fully tuning LLMs, which however will slow down inference (Hu et al., 2022). Meanwhile, the use of visual tokens greatly exacerbates the already high computation of MLLMs (Liu et al., 2023a; Zhang et al., 2023). In this paper, we focus on both parameter and computation efficient tuning for MLLMs.
2.2 Parameter and Computation Efficient Tuning
To handle the ever-increasing parameters of large-scale pre-training models, parameter-efficient transfer learning (PETL) is proposed to save the training and storage costs of LLMs (Brown et al., 2020; Li & Liang, 2021b; Hu et al., 2022). The principle of PETL is to transfer the pre-trained models to downstream tasks with only a small number of trainable parameters (Brown et al., 2020; Li & Liang, 2021b), and its methodologies can be roughly divided into three main categories, i.e. prompt tuning (Jia et al., 2022b; Zhou et al., 2022), adapter (Karimi Mahabadi et al., 2021; Gao et al., 2023a) and the re-parameterized methods (Hu et al., 2022; Luo et al., 2023a). Very recently, Wu et al. (2023b) propose a new learning task called parameter and computation efficient transfer learning (PCETL), which has an additional requirement for model inference based on the settings of PETL. And a strong baseline called Dynamic Architecture Skipping (DAS) is also proposed to evaluates the redundancy of Transformer layers by a reinforcement learning method and replaces the most redundant layers with adapters. Motivated by (Hu et al., 2022; Wu et al., 2023b), we also focus on PCETL for MLLMs in this paper. We propose a granular exploration for MLLMs, i.e., the inner MHAs, and address the latency caused by using adapters for skip connections.
PCETL is also related but orthogonal to the network compression research (Lee et al., 2023; Shi et al., 2023; Liu et al., 2023b; Kim et al., 2021a). Similar with model quantization (Lee et al., 2023; Xiao et al., 2023; Wu et al., 2023c) and pruning (Shi et al., 2023; Nova et al., 2023; Jaiswal et al., 2023), PCETL aims to improve the model efficiency on downstream tasks. A main difference is that PCETL achieves this target companied with the PETL methods like adapter, and the original parameters of MLLMs should remain unchanged due to the expensive large-scale pre-training. In this case, the contributions of PCETL and network compression are orthogonal and can be further combined. For instance, the compact model obtained by the EAS method is full-precision and can be further accelerated using quantization methods (Dettmers et al., 2023).
3 Preliminary
We first recap the principle of parameter and computation efficient transfer learning (PCETL) (Wu et al., 2023b) for multi-modal large language models (MLLMs). Concretely, given an MLLM , which consists Transformer layers and its whole parameters denoted as , the objective of PCETL can be defined by
where are the parameters of a sub-network of MLLM and is a small number of parameters for model tuning.
From Eq.1, we can see that PCETL needs to reduce parameter cost during adaption, just like PETL (Li & Liang, 2021b; Hu et al., 2022; Sung et al., 2022b). Meanwhile, it also requires practitioners to evaluate and remove the redundant modules to speed up inference. Compared with previous model compression and acceleration tasks (Krishnamoorthi, 2018; Meng et al., 2020; Dong et al., 2021), one difference of PCETL is that the original model weights should keep unchanged, so as to facilitate more task adaptions, which is essential for the giant models that are expensively pre-trained with massive data.
Wu et al. (2023b) also propose a strong baseline called dynamic architecture skipping (DAS), which considers the entire Transformer layer for redundancy evaluation. As discussed above, the components in MLLMs often serve different roles, i.e., MHA and FFN. In this case, more detailed evaluations are beneficial for PCETL. Besides, the skip connections in DAS are aided by adapters (Houlsby et al., 2019; Sung et al., 2022b) which also incur non-negligible latency during the inference.
4 Efficient Attention Skipping
4.1 Overview
To achieve parameter and computation efficient tuning for MLLMs, we propose an Efficient Attention Skipping (EAS) approach in this paper, as illustrated in Fig. 3.
Concretely, given an MLLMs , EAS aims to remove its redundant components for downstream task adaption. Different to DAS (Wu et al., 2023b), the evaluation target of EAS is more specific, i.e., MHAs of a MLLM. Thus, the objective can be defined by
where represent a subset of MHA modules and denote all FFN modules. In practice, similar to DAS (Wu et al., 2023b), we first conduct a reinforcement learning based redundancy evaluation on MHAs for downstream tasks. Afterwards, given the redundancy scores, we determine which MHAs to skip, and replace them with an adapter-based connection.
However, adapters account for excessive inference time, as shown in Fig. 1-(a). And the widely used re-parameterized methods like LoRA (Hu et al., 2022) can approximate the QKV projections but cannot replace the entire MHA for skip connections.
To this end, we propose a novel Propagation-of-Information Adapter (PIA) to achieve zero-cost skip connections, which can propagate averaged global features for information exchange, and be reparameterized into FFNs during inference.
In this case, the objective of EAS can be formulated as
where, is the reparameterizable parameters of PIA. With PIA, we can adapt MLLM to downstream tasks while speeding up inference.
4.2 Propagation-of-Information Adapter
When the redundant MHAs are evaluated, the next key step is how to effectively skip them without additional latency. To approach this target, we propose a novel re-paremeterizable adapter for EAS, termed Propagation-of-Information Adapter (PIA).
As shown in Fig. 3-(a), PIA also adopts a bottleneck structure to scale down the hidden dimension of the inputs, akin to existing adapters (Sung et al., 2022b; Luo et al., 2023a), thereby achieving a low-rank approximation of full tuning (Hu et al., 2022; Luo et al., 2023a; Wu et al., 2023a). However, it also differs in its inner path designs.
Concretely, given the input features of the layer, denoted as . PIA first projects onto two separate low-dimensional semantic spaces, and then obtain the hidden features of PIA by
where and are the two linear projections, and refers to average pooling. has a much smaller dimension than , i.e., .
In addition to the low-dimension projection, Eq. 4 also realizes the information exchange among all input tokens via the combination with the averaged feature.
During up-sampling, we also use two separate linear projections for the hidden features, and adopt a path router for their weighted combination. Thus, the output features is obtained by
Here, and are two linear projections, and is the path routing function defined by
where , is the weight matrix, is the temperature of softmax. This routing design stems from LaVIN (Luo et al., 2023b), and can handle different types of inputs, i.e., language-only or language-vision data.
Overall, the multi-path design can still hold parameter efficiency via the low-dimensional projections, while acting the role of MHA in information exchange.
Re-parameterization. With the effective structure, PIA also needs to re-parameterize its weights into the models to further improve computation efficiency. Here, we embed PIA into FFN, and avoid the skip connections (Wu et al., 2023b) to retain network complexity.
Concretely, we first place PIA in the residual connection of FFN, as shown in Fig. 3-(b):
To re-parameterize PIA into the nearby projection weights of FFN (Hu et al., 2022; Luo et al., 2022), we still need to convert PIA into one linear layer:
More specifically, we aim to simplify the definition of PIA in Eq. 4 and 5 into a two-layer linear MLP:
where and are the new projection matrices for down- and up-sampling, respectively, and , are biases.
In particular, the down-sampling for the hidden state is expected to be
However, in Eq.4, the down-sampling uses a two-path design for information exchange. In this case, we obtain and via
where and are the projection matrix and bias vector of in Eq. 4. Via Eq.11, we can merge the two-path design in one linear projection.
In practice, is frozen after the first decoding step of MLLMs regardless of the change of . This design can enable PIA to achieve information exchange during inference without hindering re-parameterization.
Similar to the down-sampling, the up-sampling, i.e. Eq.5, is redefined by
Here, and are the routing weights in Eq. 5. , are projection matrices of and in Eq.5, and are bias vectors.
After obtaining the merged and , we can then transform PIA into one linear layer via
Here, is the approximated low-rank weight matrix, and it then can be re-parameterized into the nearby weight matrices of FFN, like (Hu et al., 2022; Luo et al., 2023a).
Note that, the average feature in Eq. 4, i.e. , and the path router are example-dependent. In this case, the re-parameterization of PIA will be executed after the first decoding step of MLLMs. Considering the fact that most MLLMs often needs to decode a long sequence (Luo et al., 2023b; Zhang et al., 2023), PIA can still save massive computation during inference.
| Method | Updated Parameters | Inference Time333Tested on ScienceQA with one A100 GPU. | Subject | Context Modality | Grade | Average | |||||
| NAT | SOC | LAN | TXT | IMG | NO | G1-6 | G7-12 | ||||
| Fully tuned methods | |||||||||||
| LLaVA-7B (Liu et al., 2023a) | 100.00% | 2.19s | 90.36 | 95.95 | 88.00 | 89.49 | 88.00 | 90.66 | 90.93 | 90.90 | 90.92 |
| Parameter-efficient methods | |||||||||||
| LLaMA-Adapter (Zhang et al., 2023) | 0.03% | 4.35s | 84.37 | 88.30 | 84.36 | 83.72 | 80.32 | 86.90 | 85.83 | 84.05 | 85.19 |
| LaVIN-7B (Luo et al., 2023b) | 0.06% | 3.70s | 89.25 | 94.94 | 85.24 | 88.51 | 87.46 | 88.08 | 90.16 | 88.07 | 89.41 |
| LaVIN-7B-LoRA (Hu et al., 2022) | 0.08% | 2.22s | 84.06 | 77.95 | 85.73 | 82.84 | 74.12 | 88.64 | 83.77 | 82.20 | 83.21 |
| Parameter and computation efficient methods | |||||||||||
| DAS-7B (Wu et al., 2023b) | 0.07% | 3.23s | 88.99 | 94.60 | 85.09 | 87.88 | 86.51 | 88.36 | 89.72 | 88.13 | 89.15 |
| EAS-7B (Ours) | 0.07% | 2.13s | 89.48 | 93.93 | 87.00 | 88.32 | 87.36 | 89.76 | 90.20 | 88.99 | 89.77 |
| EAS-7B (Ours) | 0.11% | 1.69s | 89.52 | 95.61 | 86.36 | 88.71 | 88.25 | 88.85 | 91.26 | 87.67 | 89.98 |
4.3 Attention Redundancy Evaluation
In EAS, we also adopts a -arm bandit based algorithm to automatically evaluate the redundancy of MHAs in MLLMs for downstream task, similar with Wu et al. (2023b).
Concretely, given a MLLM, we first randomly sample and tune the subnetworks with PIAs as adapters for a few epochs. Afterwards, we initialize a numerical action vector, denoted as , and then we keep training and testing the MLLM with PIAs. In each evaluation step, we skip MHA modules according to , based on which the action policy at the step is obtained by
where is the uniform distribution between and . In this way, the effective modules are more likely to obtain a higher preference score. And then we skip MHA modules with the lowest preference scores:
Here, is the index of skipped MHA modules. Thus, the MHA with a greater is more likely to be kept.
Then, we update the action preference of each MHA through comparisons. For each step, we sample distinct subnetworks, i.e, , according to Eq.14 and Eq.15. Based on the loss values , the reward of a subnetwork is defined as , which is proportional to subnetwork performance.
According to the rewards, the action preference of MHA modules can be updated by
When the reward of the sampled subnetwork is larger than the average one , it suggests that the skipped MHA modules are less important. The change of their action preference in this step can be ontained by . To smooth optimization, we introduce to control the learning rate.
Via a large number of trails, most subnetworks can be well evaluated. Finally, according to the action preference , which is used to reflect the redundancy, we can select that the most redundant MHA modules to skip, thereby reducing the computation complexity of MLLMs.
| Method | Updated Parameters | VQA | NLVR | Flickr30K | Average | ||||
| test-dev | +FLOPs | test-P | +FLOPs | IR/TR R@1 | +FLOPs | Acc. | +FLOPs | ||
| Full Tuning | 323.31M | 77.43 | +0.00 | 83.05 | +0.00 | 82.22/94.30 | +0.00 | 84.25 | +0.00 |
| Classifier Only | - | 69.93 | +0.00 | 73.23 | +0.00 | 78.80/89.00 | +0.00 | 77.74 | +0.00 |
| Shallow Prompt (Li & Liang, 2021a) | 0.30M | 68.51 | +28.71G | 65.69 | +26.84G | 74.20/88.60 | +28.71G | 74.25 | +28.71G |
| Deep Prompt (Jia et al., 2022a) | 1.84M | 70.78 | +6.53G | 72.64 | +5.59G | 78.84/89.40 | +6.53G | 77.92 | +6.53G |
| LoRA (Hu et al., 2022) | 0.29M | 74.00 | +0.00 | 78.82 | +0.00 | 79.86/92.60 | +0.00 | 81.32 | +0.00 |
| Adapter (Sung et al., 2022a) | 5.34M | 74.70 | +1.64G | 79.93 | +1.38G | 80.38/91.90 | +1.64G | 81.73 | +1.64G |
| Scaled PA (He et al., 2022) | 3.59M | 75.11 | +1.12G | 80.38 | +0.66G | 80.40/93.20 | +1.12G | 82.27 | +1.12G |
| DAS (Wu et al., 2023b) | 5.34M | 74.80 | -11.16G | 80.11 | -5.13G | 80.12/91.80 | -11.16G | 81.71 | -9.15G |
| EAS(Ours) | 4.07M | 74.85 | -10.65G | 80.13 | -7.38G | 80.38/92.20 | -9.12G | 81.89 | -9.05G |
| EAS(Ours) | 4.07M | 74.82 | -14.53G | 80.08 | -10.19G | 79.78/90.40 | -11.90G | 81.27 | -12.21G |
5 Experiment
5.1 Datasets and Metrics
ScienceQA (Lu et al., 2022) is a multi-modal dataset for science question answering. This dataset is divided into three subsets, i.e., train, val, and test, which have , , and examples, respectively. The questions can be categorized into Natural Science (NAT), Social Science (SOC), Language Science (LAN), Text Context (TXT), Image Context (IMG), No Context (NO), Grades 1-6 (G1-6) and Grades 7-12 (G7-12), respectively. VQA2.0 (Goyal et al., 2017a) involves images, each accompanied by at least three open-ended questions. We report the test-dev results from the online evaluation 444https://eval.ai/web/challenges/challenge-page/830/overview. NLVR (Suhr et al., 2019a) is tailored for classifying triplets consisting of two images and a question, and it has examples of human-written sentences grounded with pairs of images. Flickr30K (Plummer et al., 2017) is a cross-modal retrieval dataset, comprising images collected from Flickr, each image has five reference sentences provided by human.
5.2 Implementation Details
The base model employed in our study is LaVIN (Luo et al., 2023b), a recently introduced MLLM. Following its default settings, we use LLaMA-7B (Touvron et al., 2023) as the main network, and adopt ViT-L (Dosovitskiy et al., 2021) as the image encoder. In terms of PIA, we set the hidden dimension to 32. Besides, we employ PIAs for parameter-efficient adaptations, similar to (Luo et al., 2023b; Zhang et al., 2023). These PIAs have a hidden dimension of 8, and only one path for adaptation is used, i.e., in Eq. 4. We insert these PIAs before each attention module. For redundancy evaluation, we first train the subnetworks with randomly skipped modules for 5 epochs. Then we introduce redundancy evaluation in the next 5 epochs, where 3 subnetworks are sampled every 10 training steps. In addition, we also validate the proposed EAS on a conventional VL pre-trained model named METER (Dou et al., 2022). Its attention modules in a co-attention layer are considered as independent potential skipped components. The PIAs for skipping attention modules are configured with a hidden dimension of 144, with the remaining settings following the defaults in DAS (Wu et al., 2023b).
| Candidate Modules | Number of Skipped modules | Updated Parameters | Inference Time | Average Performance |
| MHA | 4 | 0.08% | 2.08s | 89.44 |
| 8 | 0.10% | 1.91s | 89.79 | |
| 12 | 0.11% | 1.69s | 89.98 | |
| FFN | 4 | 0.08% | 2.26s | 88.30 |
| 8 | 0.11% | 2.18s | 84.96 | |
| 12 | 0.13% | 2.11s | 82.13 | |
| MHA or FFN | 4 | 0.08% | 2.13s | 88.40 |
| 8 | 0.10% | 2.09s | 89.18 | |
| 12 | 0.12% | 1.78s | 88.63 |
5.3 Experimental Results
5.3.1 Quantitative Analysis
Comparison with DAS. We first compare our EAS with DAS (Wu et al., 2023b) on LaVIN and METER on four VL benchmarks, of which results are given in Tab. 1 and Tab. 2. We first observe that both DAS and EAS can achieve competitive performance, while retaining parameter and computation efficiency. Speciffically, in Tab. 1, DAS can skip Transformer layers of LaVIN for better efficiency while retraining comparable performance, i.e accuracy while efficiency. Meanwhile, EAS-7B can improve the average accuracy and inference speed by and under the same parameter scale. With a little extra parameter cost, our EAS can skip MHAs while achieving better performance than both DAS and LaVIN. Notably, compared with LaVIN and DAS, EAS-7B can improve the inference speeds by and times, respectively. In Tab.2, we compare to DAS on three classic VL benchmarks, i.e. VQA, NLVR and Flickr30K. Compared with DAS, our EAS can update fewer parameters and improve inference by with only performance drop on average. Overall, these results well validate EAS towards the target of PCETL for MLLMs.
Comparison with the state-of-the-arts. In Tab. 1 and Tab. 2, we also compare EAS with more recent advancements of MLLMs and PETL methods. We can first observe that PETL methods achieve competitive or even better performance than the fully tuned methods, while retaining a very small number of parameters to update. However, the use of adapters also obviously slows down their inference speeds. For instance, the inference of LLaMA-Adapter in Tab. 1 consumes more than seconds for each example. In contrast, LoRA does not incur any additional overhead during inference, thus only taking seconds per example. Compared with these approaches, the results of EAS are more significant in both performance and efficiency. EAS-7B improve performance of LoRA while improve inference speed up to on ScienceQA. Meanwhile, in Tab. 2, we can observe that, EAS achieves performance of Scaled PA while decreasing FLOPs. Overall, these results well validate the effectiveness of EAS towards the target of PCETL for MLLMs.
Ablation Study. To examine the different designs of PIA, we further conduct extensive ablations in Tab.3-5. In Tab.3, we first ablate the choice of evaluation candidates for EAS on ScienceQA, including “MHA”, “FFN” and “MHA or FFN”. From this table, we can first see that skipping MHA is the best choice among three candidates. Notably, EAS can skip up to 12 MHAs with even better performance, while its efficiency is also much better than the others. In terms of FFN, its removal leads to slower inference speed than that of MHA. More importantly, when dropping more than 8 FFNs, the performance declines significantly, e.g., by skipping 12 FFNs. This result well confirms our argument about the roles of MHA and FFN in MLLMs. In terms of “MHA or FFN”, this candidate is suboptimal for EAS, which can obtain a good trade-off between performance and efficiency, but it is still worse than “MHA”. Overall, these results well validate the motivation of EAS, i.e., not all attention is needed for MLLMs.
Tab. 4 ablates the designs of PIA. Here, “Base” refers to using only one path in the downsampling, i.e., only in Eq. 4. We can see that without the two-path design for information exchange, its performance drops greatly, i.e., than “+information exchange”. In order to realize re-parameterization, we freeze in Eq. 11, i.e., frozen bias term in Tab. 4. It can be seen that this setting barely affects performance, i.e., . In this case, EAS can conduct re-parameterization without declining performance, while improving inference speed greatly. Overall, Tab. 4 well confirms the design of information exchange and validates the re-parameterization in PIA.
| Methods | Updated Parameters | Inference Time | Average Performance |
| LaVIN-7B | 0.06% | 3.70s | 89.41 |
| EAS-7B | 0.08% | 2.08s | 89.44 |
| EAS-7B | 0.10% | 1.92s | 89.79 |
| EAS-7B | 0.11% | 1.69s | 89.98 |
| EAS-7B | 0.13% | 1.56s | 88.42 |
In Tab. 5, we report the results of skipping different numbers of MHAs by EAS. The first observation is that skipping appropriate MHA modules has little impact on performance, e.g. skipping up to 16 MHA modules only has about performance drops, strongly suggesting that PIA is a good substitute for MHA. Notably, skipping a certain numbers of MHAs can achieve even better performance than the default LaVIN, e.g. -12 layers with accuracy. It might be due to the great redundancy of MLLMs on ScienceQA, which also suggests that PIA can help MLLMs learn better patterns from VL data. Again, these results confirm the effectiveness of EAS towards PCETL.
5.3.2 Qualitative Analysis
Examples of Science Question Answering. In Fig. 4, we visualize the examples of EAS-7B and LaVIN-7B for both language-only (left) and image-language questions (right). From these examples, we can see that both EAS-7B and LaVIN-7B correctly answer the question. Meanwhile, they all inherit the strong language ability of LLaMA (Touvron et al., 2023), and can explain the answers fluently. Notably, for the visual question (right), EAS-7B has better and more detailed explanations for the answer, e.g., “The colony’s name comes from the fact that it was founded by people from Hampshire, a county in England”. In contrast, the response of LaVIN is logically incorrect even through its answer is right. Overall, these results confirm again the merits of EAS for the efficient adaption of MLLMs.
6 Conclusion
In this paper, we propose a novel method for parameter and computation efficient tuning of Multi-modal Large Language Models (MLLMs), named by Efficient Attention Skipping. Concretely, we first reveal that not all MHAs are necessary for the efficient adaption of MLLMs, based on which EAS adopts a granular redundancy evaluation scheme. Meanwhile, to avoid the additional computation caused by the adapter-based skip connections, EAS is also equipped with a novel Propagation-of-Information Adapter (PIA), which can not only keep parameter efficiency but also can be re-parameterized into the model without extra latency. To validate EAS, we apply it to a recent MLLM called LaVIN, and conduct extensive experiments on the multi-modal benchmark ScienceQA. The experimental results show that EAS can achieve better performance than LaVIN while speeding up inference by up to 2.18 times.
7 Brader Impact
This paper address the parameter and computation efficiency of MLLMs through efficient attention skipping (EAS), which will further advance the development of MLLMs in the machine learning community. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Influence of hidden dimension
| Methods | Hidden Dimension | Updated Parameters | Inference Time | Average Performance |
| EAS-7B | 8 | 0.04% | 1.89s | 87.64 |
| 16 | 0.05% | 1.91s | 87.95 | |
| 32 | 0.08% | 1.92s | 89.79 | |
| 64 | 0.22% | 1.94s | 90.07 |
Figure 5:
Comparison between EAS and DAS on ScienceQA.
Tab. 6 reports the impact of different trainable parameter sizes on EAS via setting the hidden dimension. From the table, we can observe that there is a significant performance increase with more updated parameters, e.g. the performance is improved by with updated parameters more. This result also suggests that replacing redundant MHAs still needs a certain model capacity to accommodate. Even so, the optimal solution, i.e., skipping 12 MHAs, still consumes a very small proportion of parameters, i.e., , showing high parameter efficiency of EAS.
Appendix B Comparison with DAS
We first compare our EAS with the strong baseline DAS under different experimental settings in Fig. 5. The first observation from Fig. 5 is that DAS can skip certain layers of LaVIN with very limited performance drops, e.g., skipping 2-4 layers. Its efficiency gains become more obvious when skipping more than 6 Transformer layers, while the performance also decreases to a certain extent. Compared with DAS, EAS has obvious merits in both performance and efficiency. For instance, with the same number of skipped modules, EAS is consistently faster than DAS by to . By skipping 8 MHAs, EAS can obtain more obvious performance gains than the default LaVIN, i.e., 89.79 v.s. 89.41, while the inference speed is about times faster. Overall, these results well validate the motivation and designs of EAS towards PCETL of MLLMs.
| Method | Updated Parameters | Training Memory | Inference Time | Subject | Context Modality | Grade | Average | |||||
| NAT | SOC | LAN | TXT | IMG | NO | G1-6 | G7-12 | |||||
| LaVIN-7B (Luo et al., 2023b) | 3.8M | 32.32G | 3.70s | 89.25 | 94.94 | 85.24 | 88.51 | 87.46 | 88.08 | 90.16 | 88.07 | 89.41 |
| DAS-7B (Wu et al., 2023b) | 4.2M | 32.44G | 3.44s | 88.68 | 94.94 | 86.45 | 88.03 | 86.81 | 88.92 | 90.20 | 88.00 | 89.41 |
| DAS-7B (Wu et al., 2023b) | 4.6M | 31.67G | 3.23s | 88.99 | 94.60 | 85.09 | 87.88 | 86.51 | 88.36 | 89.72 | 88.13 | 89.15 |
| DAS-7B (Wu et al., 2023b) | 5.0M | 30.91G | 3.06s | 87.30 | 93.36 | 82.36 | 86.12 | 85.97 | 85.71 | 88.18 | 85.70 | 87.29 |
| DAS-7B (Wu et al., 2023b) | 5.4M | 30.05G | 2.94s | 86.50 | 89.31 | 84.27 | 85.63 | 83.74 | 86.55 | 87.41 | 84.90 | 86.51 |
| EAS-7B(Ours) | 4.5M | 32.82G | 2.13s | 88.81 | 95.16 | 85.36 | 87.29 | 86.32 | 88.99 | 90.35 | 87.28 | 89.25 |
| EAS-7B(Ours) | 5.1M | 32.50G | 2.08s | 88.81 | 94.83 | 86.36 | 87.68 | 86.81 | 89.06 | 90.42 | 87.67 | 89.44 |
| EAS-7B(Ours) | 5.7M | 32.05G | 1.99s | 88.99 | 94.15 | 85.36 | 87.88 | 87.46 | 88.92 | 90.68 | 86.68 | 89.35 |
| EAS-7B(Ours) | 6.4M | 31.63G | 1.92s | 90.05 | 94.04 | 85.82 | 89.25 | 87.56 | 88.78 | 90.68 | 88.20 | 89.79 |
| EAS-7B(Ours) | 7.7M | 31.10G | 1.69s | 89.52 | 95.61 | 86.36 | 88.71 | 88.25 | 88.85 | 91.26 | 87.67 | 89.98 |
Appendix C The Detailed Results
We further report the detailed experimental results of DAS and EAS for different numbers of skipped modules. From Tab.7, we observe that the performance of DAS drops faster than the proposed EAS, especially in the TXT benchmark, which states that the skipping of FFN will lead to a weakness in language ability. Meanwhile, we notice that the improvement of inference efficiency in DAS is much slower than in EAS. Overall, the results further prove the effectiveness of our EAS.
Appendix D The detailed skipping results
We first visualize the accelerated MLLM searched by DAS (Wu et al., 2023b) and our EAS. From Fig. 6, we can observe that the layers skipped by both DAS and EAS tend to be deep layers. While the layers that are considered to have redundancy in EAS are deeper than those in DAS. Furthermore, we can find out that the skipped modules are concentrated behind the middle of MLLM, and gradually expand forward as the number of skipped modules increases.